RÉPUBLIQUE ET CANTON DE GENÈVE

OBSERVATOIRE DE GENÈVE

CH-1290 SAUVERNY Suisse

Dr D. Mégevand Ingénieur de recherche Tel: +41 (22) 755 26 11 Fax: +41 (22) 755 39 83

Sauverny, January 7, 2002

TEX Users Group P.O.Box 594 Providence, RI 02901 USA

Y.^{/}ref. TUG/kb  O.^{/}ref. PCT8289  C.C.P. 12-2130-4

Subject: PiCTEX Manuals

Dear Sir,

Enclosed, you will find a check of $60.00 corresponding to your invoice #902 dated June 2,1993 for 2 PiCTEX manuals.

Sincerely

Dr D. Mégevand

Encl(s): 1 check

E-Mail:  megevand@obs.unige.ch